What is a better buy? A square pizza measuring 8 inches that costs $10 or a round pizza with a 9-inch diameter that also costs $10?
Answer by aaronwiz(69)

area of square pizza=8*8=64
area of circle pizza=4.5^2(pi)=63.61
so the square pizza will give you slightly more